Determine the (H+), OH), and pH of a solution with a pok of 12.49 at 25 °C. [OH-] = pH =
Answer #1

use:

PH = 14 - pOH

= 14 - 12.49

= 1.51

use:

pH = -log [H+]

1.51 = -log [H+]

[H+] = 3.09*10^-2 M

use:

pOH = -log [OH-]

12.49 = -log [OH-]

[OH-] = 3.236*10^-13 M

Answers:

[H+] = 3.1*10^-2

[OH-] = 3.2*10^-13

pH = 1.51